The prior probabilities for events A1 and A2 are P(A1) = 0.40 and P(A2) = 0.45. It is also known that P(A1A2) = 0. Suppose P(B | A1) = 0.20 and P(B | A2) = 0.05. If needed, round your answers to three decimal digits.

(a) Are A1 and A2 mutually exclusive?
- Select your answer -YesNoItem 1
Explain your answer.
The input in the box below will not be graded, but may be reviewed and considered by your instructor.
(b) Compute P(A1B) and P(A2B).
P(A1 B) =
P(A2B) =
(c) Compute P(B).
P(B) =
(d) Apply Bayes’ theorem to compute P(A1 | B) and P(A2 | B).
P(A1 | B) =
P(A2 | B) =
Answer #1

a)Yes since P(A1 n A2) =0 , therefore A1 and A2 are mutually exclusive

b)

P(A1 n B) =P(A1)*P(B|A1) =0.4*0.2 =0.08

P(A2 n B) =P(A2)*P(B|A2) =0.45*0.05 =0.0225

c)

P(B) =P(A1 n B)+P(A2 n B) =0.1025

d)

P(A1|B) =P(A1 n B)/P(B) =0.7805

P(A2|B) =P(A2 n B)/P(B) =0.2195